When compounds of barium are heated in a flame, green light of wavelength 554 nm is emitted. How much energy is lost when one mole of barium atoms each emit one photon of this wavelength?
1
Expert's answer
2019-05-27T06:10:28-0400

Energy of foton is:

E=h(c/l)

Where h - Planck constant

c - light speed in vacuum

l - wavelength

1 mol of Ba emit one mol of photons, then

E = h(c/l)NA5= 2.16*105 J